### v1

# StarDist: A Code Generator for Distributed Graph Algorithms

Abstract
Relational data, occurring in the real world, are often structured as graphs, which provide the logical abstraction required to make analytical derivations simpler. As graphs get larger, the irregular access patterns exhibited in most graph algorithms, hamper performance. This, along with NUMA and physical memory limits, results in scaling complexities with sequential/shared memory frameworks. StarPlat’s MPI backend abstracts away the programmatic complexity involved in designing optimal distributed graph algorithms. It provides an instrument for coding graph algorithms that scale over distributed memory. In this work, we provide an analysis-transformation framework that leverages general semantics associated with iterations involving nodes and their neighbors, within StarPlat, to aggregate communication. The framework scans for patterns that warrant re-ordering in neighborhood access patterns, aggregate communication, and avoid communication altogether with opportunistic caching in reduction constructs. We also architect an optimized bulk-reduction substrate using Open MPI’s passive Remote Memory Access (RMA) constructs. We applied our optimization logic to StarPlat’s distributed backend and outperformed d-Galois by 2.05 and DRONE by 1.44 times in Single Source Shortest Paths across several big data graphs.
I Introduction
For distributed graph algorithms to perform reasonably, the cost of communication must compensate that of page faults, critical section overheads and inefficient RAM-processor affinity (NUMA) utilization frequently noted in sequential alternatives.. On compute clusters with SAN storage and RDMA enabled network interfaces, network communication is often faster than storage[8].
IO intensive workloads on large data, give distributed algorithms a scope to outperform shared memory or sequential algorithms. The time, otherwise spent on page faults, would be utilized in computing a partial-result on another node, and the communication of the computed result would be faster than a fault. Graph algorithms are IO intensive, due to their irregular memory access patterns. Even for reasonably sized graphs, the cache miss rate is considerable[9].
Figure 1: Left: StarPlat Right: DSL code sample.
codeASTmpigpucudahipsycloclompnlyz
forall v in g.nodes(){
              forall nbr in g.neighbors(v){
                nbr.color = <Sum<Max(v.color,
                            nbr.color)>, 1>;
Dividing the work among different processes reduces the amount of memory managed per process. This allows different processes, assigned to different cores, to use a different L1(private) cache, the data ownership ensures most of the caching activity is restricted to a subgraph. StarPlat is a DSL compiler, tailored toward scaling graph algorithms, with a code once, run anywhere principle. It allows generation of multiple targets based on the programmer’s requirement. This restricts us to a unified set of semantics, which is of greater importance than one target code’s efficiency.
Performance aspects unique to the distributed backend are world involvement in point to point synchronization and overheads in initiating communication which cause notable slowdown. This backend also has a per participant, multiple vertices problem, i.e, vertex centric view is not feasible, unlike it’s counterparts on the shared memory backends, where a thread per vertex is launched.
In this paper, we introduce the following concepts that have given us an edge in performance :
  1. 1.
    A bulk-reduction substrate backed by passive MPI RMA calls in junction with a distributed work list. This eliminates the need for receiving ranks to be aware of ranks and sizes of incoming packets and also exploits RDMA capabilities of underlying hardware[2].
  2. 2.
    We define reduction-exclusive statements that provide a stub for achieving optimal communication patterns. 1b is an example where both iterations are reduction-exclusive. Reduction-exclusive statements allow delayed updates enabling our work to perform on par with Gluon[1] used in Galois [3] which utilize synchronization passes for synchronizing master and mirror nodes.
  3. 3.
    We add an analysis-transformation component, specific for StarPlat’s distributed backend. Profiling of various target codes revealed many bottlenecks that can be levied by basic backend analysis. These analyses focus on bypassing bulky function calls and aggregating communication requirements to maximize CPU and bandwidth utilization, and optimizing neighborhood traversals the traversal order does not matter.
  4. 4.
    We reduce communication overhead by opportunistic caching of foreign data within reduction-exclusive statements.
II Related Works
Most of the prior works on distributed graph algorithm frameworks are library backed, requiring users to arrange a sequence of API calls to map out their algorithm.
StarPlat provides a unified high performant, domain specific, compiled language with an Abstract Syntax Tree representing semantics capturing graph entities and operations. However, it struggles with generating optimized code for distributed graph algorithms, which our work solves, allowing users to code simple graph algorithms that generate optimal distributed graph algorithms.
Distributed Galois[1] provides a framework leveraging gluon-async[1] that can be used to write peer-to-peer distributed graph algorithms. Gluon Async[1] provides a bulk asynchronous parallel model which provides mechanisms to bypass global locks on distributed memory and allows passive updates across participants through proxy nodes.
Each vertex in this paradigm may have multiple proxy nodes, out of which, one is elected master, responsible for holding the canonical value, while the others are regarded as mirror nodes.
Gemini [4] framework adapts to runtime and provides an effective edge centric graph partitioning scheme. It facilitates writing graph algorithms both in shared and distributed memory situations. Much like gluon async [1] in distributed Galois [3], updates are propagated iteratively until convergence.
DRONE[2] explores a subgraph boundary synchronization scheme over a vertex-cut based partitioning, designed similar to d-Galois but with the additional clause of only boundary nodes having master and mirror nodes and a master-slave architecture where the master node responsible for synchronization and book-keeping. DRONE [2] claims better performance over gluon[1] and gemini [4]
In all three of the above frameworks [1] [4] [2], the propagation of updates to mirror nodes and their synchronization happens in pulses, as defined in Goldberg and Tarjan’s parallel implementation of push relabel[8]. The convergence paradigm represents an inherent universality as it works across shared memory multi-core, distributed memory, as well as GPUs. However, for distributed Graph algorithms, every pulse is expensive and bulkier pulses happening less frequently achieving similar results is a faster directive.
Poly[10] The exploitation of reduction operations in polyhedral analysis are detailed in [10] where multiple possible reduction operations were evaluated to perform 2.21 times faster than control. Semantic knowledge of delay-able statements present in reduction operators allowed scheduling of multi-level iterations without both memory and control barrier delays.
Cagra[9] The irregular memory access patterns with graph algorithms and their effects when scaled are studied in Cagra[9], that attempts to bring down the cache miss rate by vertex relabeling. Cagra [9] does a best effort service to fit the neighborhood of a vertex into the L1/L2 cache of the CPU for shared memory frameworks. Distributed memory, especially MPI, would have one process affined to a core, and the vertices managed by a core would be a dividend of the total vertices present in the graph, thus achieving a similar effect with cache misses, in a natural fashion.
StarPlat[5] a versatile graph DSL provides a distributed backend for generating graph algorithms in openMPI, a more detailed explanation of it is within section III. Whereas inclusion of RMA to StarPlat in [6] gave flexibility to code generation, the performance regressed multiple times, small and frequent RMA updates are expensive due to RMA read/update cycle initiation and termination overheads.
The StarPlat compiler’s[6] distributed backend utilizes a peer-to-peer graph data structure in its header library and generates the openMPI code from the DSL code using StarPlat’s graph AST. In this section we describe essential structures in the StarPlat’s distributed backend.
sg1sg2sg3
(a) graph data structure
funcsigsigblkwhlforifcondbodydeclasgnexpr
(b) abstract syntax subtree
Figure 2: Background
  • Graph Data Structure: The graph partitions, as shown in Figure 2a, in Starplat[5] do not contain ghost or mirror nodes used in gluon-async [1] or DRONE [2]. Updates are directly and instantaneously propagated, and the pulses used to propagate updates in [1], [2] and [4] are not implemented so that the programmer using StarPlat does writes their code without having any awareness of parallel programming. This results in frequent updates that are smaller in size (usually 4 Byte integers).
    The overhead of initiating and terminating RMA cycles for frequent, small transmissions, is not handled by openMPI[7]. Therefore, the performance dipped for static Single Source Shortest Paths from [5] to [6] from 3411s to 6331.941s, an increase of 1.856 times.
  • StarPlat’s AST: StarPlat uses its own abstract syntax tree as a form of MLIR. The traversal provides statements. Each statement has a type parameter which it is type casted into during mpi analysis and code generation. This type casting enables traversal of the tree. A sub-tree consisting the ast is shown in Figure2b .
  • Reduction Operations: Reduction operations for StarPlat MPI provide a different set of semantics. They allow aggregating necessary updates. The operations are queued in a reduction queue maintained in the graph structure described in subsection A. Once the queuing is done, a sync_reduction is called which starts the actual updates.
III Motivation
The StarPlat compiler’s distributed backend prefers generality in code generation over performance. By profiling generated code for Single Source Shortest Paths using gprof [11], perf and hpctoolkit[12], we deduced function calls that were utilizing high number of CPU cycles. We logged times of these functions and concluded the following phases to be expensive:
  • Reduction: The reduction operation provides an opportunity to propagate in bulk after queuing updates for an iteration. Propagating the updates in a synchronize reduction step takes a good portion of time in graphs with high average vertex degree as is evident from 3c.
  • get calls on local nodes: Get calls on MPI RMA require locking down the window in order to achieve mutual exclusion, so that another entity cannot read the same node’s value. In Figure 3d, the amount of time spent in get calls is almost as much as the time spent in synchronizing reduction, and is dependent heavily upon the amount average vertex degree of the graph.
  • get calls on foreign nodes: Get calls on MPI RMA to foreign nodes have the undeniable overhead of reading in the vertex property from a different process’ context. However, between subsequent get calls, these calls can be avoided even if the value did not change.
  • get_edges : Profiling graph traversal highlighted a reasonable amount of time being spent in acquiring the edges. The get_edge() function calls for neighborhood traversal incurs higher cost for graphs with large diameters, but is also high in graphs with high average vertex Degree, [figure 3a]
Together, these three activities take up a huge chunk of the total execution time as shown in figure 8b. We also observe that using the reduction construct in the DSL alone, could provide domain knowledge that could reduce the get calls, as well as, reduce the size of the reduction queue.
(a) time spent in edge query in neighborhood traversal.
(b) nbr traversal, sync red and get calls together
(c) Time spent in synchronizing reduction.
(d) Time spent in get calls
Figure 3: Profiling observations
To this end, we formalize the below definitions which help implement safety analysis, which when passed, we apply transformations to optimize generated code.
  • Definition 1.
    A statement S, AST traversal of which can lead to exactly one reduction statement R which is reading/updating elements in E, and E is neither accessed nor updated outside R can be defined reduction-exclusive
  • Definition 2.
    A property P is defined opportunistic cache safe if P is not updated within a reduction-exclusive statement.
  • Definition 3.
    We define an inter process synchronization following a period of intra process computation as a pulse.
  • Lemma 1.
    A statement S within a reduction-exclusive statement is also a reduction-exclusive statement.
The definition of a pulse 3 is an extended, more abstract form of what appears in [8], where delayed and aggregated communication was found beneficial for distributed algorithms.
The lemma 1, allows us to aggregate several pulses 3 and achieve bulkier and less frequent communication, which is ideal.
IV The Backend Analyzer
We introduce an analyzer framework for the distributed backend for StarPlat to add specificity toward commonly occurring patterns that would benefit from avoiding taking global locks and reduced communication.
Neighborhood traversal: Traversing the neighbors of a set of vertices is a staple pattern for most graph algorithms, and in StarPlat, any process can access the edges and vertices in the neighborhood of a vertex belong to any process in the world.
To adhere to the peer-to-peer and no mirror node ideology, StarPlat initiates communication with owner nodes and transfers the neighborhood information on demand. We noticed through gprof[11], that get_edge calls within neighborhood traversals were taking up a high number of CPU cycles.
(a) DSL input.
# DSL code
forall v in g.nodes(){
forall nbr in g.neighbors(v){
Edge e = g.get_edge(v,nbr);
(b) Current generated code.
forall v in g.nodes(){
forall nbr in g.neighbors(v){
Edge e = g.get_edge(v,nbr);
(c) optimal generated code.
forall v in g.nodes(){
int _t1 = 0;
forall nbr in g.neighbors(v){
Edge e = g.get_edge_i(v,_t1);
nbr = g.get_edge_other(v,e);
_t1++;
Figure 4: Neighborhood traversal proof of correctness.
A particular edge access involved searching through the neighborhood edges. In a pattern like in 4 the search would add O(n) traversal time for every vertex whose neighborhood required an edge. This combined with the fact that the forall node traversal in StarPlat also indicates parallel traversal and therefore must be executed by all processes, gives us semantic information that nodes v are local and their neighborhood’s CSR must be local as well. StarPlat’s MPI header stores edges consecutively, therefore, if an edge is accessed closed to a previously accessed edge, the cache misses are reduced drastically. Therefore, we rearrange the neighborhood traversal, as shown in Figure 4, to match the order present in the adjacency array of the local CSR, thereby eliminating the need for searching for an edge entirely.
Aggregating reduction pulses: A common pattern of calculating and propagating updates of various vertex and edge properties in a distributed setup is through pulses in definition 3, as seen in SSSP, Maximum flow and CC. Within the pulse, the what-to-update is decided and across two pulses the updates are propagated.
By leveraging lemma 1, we extend the reduction statement in StarPlat to form pulses out of reduction-exclusive statements. A section on bulk_reduction provides enhancements we introduce to leverage reductions even more. The reduction operations being commutative and associative, does not require a maintenance of ordering of operations.
(a) DSL input.
# DSL code
while(!g.reduction_fronter())
forall(v in g.reduction_frontier())
<nfs.x> = <Min (w.x, v.x)>;
(b) Current generated code.
while(!g.reduction_frontier())
for(int v : g.reduction_frontier())
g.add_red(nbr.get(x), v.get(x));
g.sync_reduction();
(c) optimal generated code.
while(!finished)
while(!g.reduction_frontier())
for(int v : g.reduction_frontier())
g.add_red(nbr.get(x), v.get(x));
g.sync_reduction();
if(!g.reduction_frontier())
finished = true;
MPI_Allreduce(&finished, &finished, MPI_LOR);
Figure 5: Aggressive Aggregation into pulses.
MPI_Get bypass: The get calls for RMA are quite expensive even for vertices local to a process.
  • The RMA phase needs to be initiated.
  • The window needs to be locked.
  • The data must be transferred, and window must be unlocked.
  • RMA phase needs to be terminated.
  • The stack depth for an MPI_Get call is huge.
OpenMPI[7] does not internally optimize subsequent RMA calls by keeping RMA phases active. Keeping the RMA Window under a shared lock also does not optimize the get calls. We use lemma 1 and our analysis for determining reduction-exclusive statements as in definition 1 to argue that get calls within statements deemed reduction-exclusive can safely de-reference the local memory of an RMA window, as updates to the data undergoing reduction are not going to be propagated until the end of the pulse. We show our analysis in Figure 6
(a) DSL input.
# DSL code
while(!g.reduction_fronter())
forall(v in g.reduction_frontier())
<nfs.x> = <Min (w.x, v.x)>;
(b) Current generated code.
while(!g.reduction_frontier())
for(int v : g.reduction_frontier())
g.add_red(nbr.get(x), v.get(x));
g.sync_reduction();
(c) optimal generated code.
while(!g.reduction_frontier())
for(int v : g.reduction_frontier())
{ if(v is local){
a = x.localdata[local(v)];
b = x.localdata[local(w)];
if(a>b)
x.localdata[local(v)] = y;
}else
g.add_red(nbr.get(x), v.get(y));
g.sync_reduction();
Figure 6: get call optimization in reduction operations into pulses.
opportunistic caching: In reduction exclusive statements [definition 1], we note that MPI_Get calls going toward a vertex residing in another process’ address space need not be brought into the local memory twice. The vertex property values will not change until the end of the pulse and therefore maintaining and checking for a dirty bit globally is also not necessary.
(a) DSL input.
# DSL code
forall v in g.nodes(){
forall nbr in g.neighbors(v){
Edge e = g.get_edge(v,nbr);
<nbr.a> = <Sum (1,nbr.a)>;
(b) Current generated code.
map cache_v;
map cache_nbr;
forall v in g.nodes(){
forall nbr in g.neighbors(v){
Edge e = g.get_edge(v,nbr);
g.add_to_red(1,cache_nbr);
g.sync_reduction();
cache_v.clear();
cache_nbr.clear();
(c) optimal generated code.
map cache_v;
map cache_nbr;
forall v in g.nodes(){
forall nbr in g.neighbors(v){
Edge e = g.get_edge(v,nbr);
if(cache_nbr.find(v)
!= cache_nbr.end())
cache_nbr[nbr]=a.get_val(nbr);
g.add_to_red(1,cache_nbr);
g.sync_reduction();
cache_v.clear();
cache_nbr.clear();
Figure 7: Opportunistic caching.
V Bulk Reduction
$P0P1PN$
(a) all-to-all congestion.
$P0P1PN$
(b) bulk-reduction rma calls
Figure 8: Communication Profile before and after
graphrmagpnpep
(a) current StarPlat design.
graphrmagpnpepblkRGlobal struct singleton
(b) bulk reduction substract integrated design
Figure 9: StarPlat MPI header structure before and after
As mentioned in section IV, the pulses formed from aggregation cause the synchronization of the reduction calls to get bulkier, and from section III, the synchronize reduction operation is an expensive call. We design a bulk_reduction substrate shown in Figure 9 with the goal to:
  • Propagate updates by causing minimal communication overhead.
  • Minimize the number of updates going onto the RMA window.
  • Cache friendly reduction queue.
Reduction Queue: To obtain a cache efficient reduction queue, world size number of list of arrays where each array is of a size slightly smaller than the size of the L1 cache divided by world size was chosen as the appropriate data structure. Once an array is full, we allocate a new array at the end of the list. An stl vector in C++ has doubles its capacity and we’ve observed higher cache misses once the vector’s size starts exceeding that of the cache. An update going toward process $i$ would be populated in the $i_{th}$ list of arrays. Adding an element to the reduction queue is illustrated in algorithm 1 The choice of this data structure arose from the following observations:
  • Once an element is pushed into the reduction queue, it is not accessed until the next synchronize reduction call.
  • During a synchronize reduction call, the entire reduction queue is traversed exactly once.
Therefore, a memory block is accessed twice, once when an element is inserted into the queue and once again during synchronize reduction. An STL vector is optimized for an undefined number of accesses and its amortization works well in the general case. Each element in the array of the list would comprise global index of the property being updated, and the value with which it is to be updated.
Algorithm 1 ADD TO REDUCTION
procedure add to red(idx val)
  if buf_count<BUFFERSZ/worldsizebuf\_count<BUFFERSZ/worldsize then
   $buf\_xfer\leftarrow malloc(BUFFERSZ/worldsize)$
   $buf\_count\leftarrow buf\_count+1$
  end if
  $buf\_xfer[buf\_count][buf\_idx]\leftarrow idx$
  $buf\_idx\leftarrow buf\_idx+1$
  $buf\_xfer[buf\_count][buf\_idx]\leftarrow val$
  $buf\_idx\leftarrow buf\_idx+1$
end procedure
Short circuiting reduction: The updates to a reduction call are not needed within a reduction-exclusive statement. Therefore, if the reduction operation is monotonic, the distributive property allows us to update the values for local vertices or edges immediately as get calls to the vertices within a reduction-exclusive statement can expect the updates to be propagated without being utilized anywhere other than the reduction itself.
Synchronizing reduction: We define an RMA window with each process having a local contribution equal to the size of the individual arrays in the reduction queue list. While synchronizing, as illustrated in algorithm 1:
  • A passive shared lock is taken with the no_check_neededno\_check\_needed flag set.
  • Every process traverses the list of arrays in reduction queue, one destination at a time.
  • This way, the xmm registers in Intel Xeon are exploited for vectored moves.
  • Once the movement is completed, the processes read their regions of the RMA window, and update the respective node property.
The transfer happens in chunks until all the updates are propagated.
The legacy synchronize reduction used an all-to-all that congested the network, which is hard to scale. The RMA approach in our bulk reduction substrate is cache friendly and minimizes communication overhead by eliminating receiver awareness of incoming data.
Algorithm 2 BULK SYNC algorithm
procedure bulk synchronize(Window, graphProp) whilebulk_xfer\ bulk\_xfer
  for $rank\ \epsilon\ world$ do
   $send\leftarrow bulk\_xfer[rank]$
   $Window[rank*BUFFERSZ]\leftarrow send$
   $update\leftarrow Window[rank*BUFFERSZ]$
   for $i\leq BUFFERSIZE$ do
     $idx\leftarrow Window[rank+offset]$
     $val\leftarrow window[rank+offset+1]$
     $offset\leftarrow offset+2$
     $gProp[idx]\leftarrow reduction_{op}(gProp[idx],val)$
   end for
  end for
end procedure
VI Experimentation
Our experimental includes three nodes with Intel Xeon 32 core processors, 192 GBs of RAM. Each node has 2 NUMA regions and the three nodes are interconnected via infiniband. 32 processes per node communicating with each other stresses the network, as well as simulates a real world situation where network congestion is comon place.
VI-A DATASET
We perform our experiments on the following test suite containing a mixture of social media and road network graphs. We’ve added weights of $[0,100]$ to all the graphs.
TABLE I: Graph Dataset Characteristics
Graph Acronym Vertices Edges
(million) (million)
twitter-2010 TW 21.2 265.0
soc-sinaweibo SW 58.6 261.0
orkut OK 3.0 234.3
wikipedia-ru WK 3.3 93.3
livejournal LJ 4.8 69.0
soc-pokec PK 1.6 30.6
usaroad US 24.0 28.9
germany-osm GR 11.5 12.4
rmat876 RM 16.7 87.6
uniform-random UR 10.0 80.0
VI-B Comparison with similar arts
DRONE kept timing out in its graph partitioning phase when operated at our maximum compute capacity. Therefore, we run two sets of comparison. One with DRONE, with 2 servers, 2 NUMA each, and a total of 64 processes. We run 60 processes while allocating 64 processors to allow tcp dog threads to communicate. Both Galois and DRONE used hardware acceleration, Galois through pthreads and DRONE through CUDA. Our work focuses on the optimizing the communication pattern. Leveraging hardware acceleration for intra-node activities is fairly trivial and obscures away communication patterns that require optimization. To this end, we disable all hardware acceleration in both frameworks for a fair comparison.
For Connected Components, we use iterBFS construct in StarPlat. The header is aware of the level order topology and could simply return the number of elements in level 0 as the number of connected components. However, we feel that traversing the graph in iterBFS and using reductions truly emphasizes the improvements we have made in this work.
TABLE II: Performance comparison for $C$ across frameworks (60 Procs). All times in seconds.
Graph DRONE Galois StarPlat
TW 66.511 7.326 34.435
OK 12.827 11.729 4.754
WK 7.799 10.493 3.737
LJ 8.018 11.772 3.853
PK 4.604 9.237 1.692
US INF 6.728 52.696
GR 221.532 70.971 29.093
RM 10.269 13.267 12.232
UR 39.515 21.603 2.651
Total 660.806 163.026 145.143
TABLE III: Performance comparison for SSSP across frameworks (60 Procs). All times in seconds.
Graph DRONE Galois StarPlat
TW 34.419 6.490 0.078
SW 0.532 6.627 0.093
OK 30.646 36.831 38.117
WK 14.369 99.816 14.612
LJ 20.59 25.658 13.722
PK 9.988 19.389 4.465
GR 12.219 176.667 10.966
RM 2.814 16.475 21.645
UR 41.13 370.000 12.058
Total 166.707 424.952 115.759
Figure 10: Comparison of SSSP of StarPlat with bAnalyzer.
VII Conclusion
The backend analyzer can be expanded to detect reduction statements from operators. The reduction operator still requires the StarPlat programmer to be aware of parallel aspects of the code. A control flow graph and a data flow graph along with the semantic context provided by the Abstract Syntax Tree can abstract away the need for the Reduction operator altogether for a multitude of cases.
VIII NOTE:
The authors thank Atharva Chougule and Rohan Kumar for preliminary discussions regarding the METIS integration through vertex relabeling, and aggregation of atomic Writes. Further contributions to this work are anticipated.

### v2

@@
 For distributed graph algorithms to perform reasonably, the cost of communication must compensate that of page faults, critical section overheads and inefficient RAM-processor affinity (NUMA) utilization frequently noted in sequential alternatives..
-On compute clusters with SAN storage and RDMA enabled network interfaces, network communication is often faster than storage[8].
+On compute clusters with SAN storage and RDMA enabled network interfaces, network communication is often faster than storage.
 IO intensive workloads on large data, give distributed algorithms a scope to outperform shared memory or sequential algorithms.
@@
 Graph algorithms are IO intensive, due to their irregular memory access patterns.
-Even for reasonably sized graphs, the cache miss rate is considerable[9].
+Even for reasonably sized graphs, the cache miss rate is considerable[ref9].
 Figure 1: Left: StarPlat Right: DSL code sample.
@@
     A bulk-reduction substrate backed by passive MPI RMA calls in junction with a distributed work list.
-This eliminates the need for receiving ranks to be aware of ranks and sizes of incoming packets and also exploits RDMA capabilities of underlying hardware[2].
+This eliminates the need for receiving ranks to be aware of ranks and sizes of incoming packets and also exploits RDMA capabilities of underlying hardware[ref2].
   2. 2.
@@
 1b is an example where both iterations are reduction-exclusive.
-Reduction-exclusive statements allow delayed updates enabling our work to perform on par with Gluon[1] used in Galois [3] which utilize synchronization passes for synchronizing master and mirror nodes.
+Reduction-exclusive statements allow delayed updates enabling our work to perform on par with Gluon[ref1] used in Galois [ref3] which utilize synchronization passes for synchronizing master and mirror nodes.
   3. 3.
@@
 However, it struggles with generating optimized code for distributed graph algorithms, which our work solves, allowing users to code simple graph algorithms that generate optimal distributed graph algorithms.
-Distributed Galois[1] provides a framework leveraging gluon-async[1] that can be used to write peer-to-peer distributed graph algorithms.
-Gluon Async[1] provides a bulk asynchronous parallel model which provides mechanisms to bypass global locks on distributed memory and allows passive updates across participants through proxy nodes.
+Distributed Galois[ref1] provides a framework leveraging gluon-async[ref1] that can be used to write peer-to-peer distributed graph algorithms.
+Gluon Async[ref1] provides a bulk asynchronous parallel model which provides mechanisms to bypass global locks on distributed memory and allows passive updates across participants through proxy nodes.
 Each vertex in this paradigm may have multiple proxy nodes, out of which, one is elected master, responsible for holding the canonical value, while the others are regarded as mirror nodes.
-Gemini [4] framework adapts to runtime and provides an effective edge centric graph partitioning scheme.
+Gemini [ref4] framework adapts to runtime and provides an effective edge centric graph partitioning scheme.
 It facilitates writing graph algorithms both in shared and distributed memory situations.
-Much like gluon async [1] in distributed Galois [3], updates are propagated iteratively until convergence.
-DRONE[2] explores a subgraph boundary synchronization scheme over a vertex-cut based partitioning, designed similar to d-Galois but with the additional clause of only boundary nodes having master and mirror nodes and a master-slave architecture where the master node responsible for synchronization and book-keeping.
-DRONE [2] claims better performance over gluon[1] and gemini [4]
-In all three of the above frameworks [1] [4] [2], the propagation of updates to mirror nodes and their synchronization happens in pulses, as defined in Goldberg and Tarjan’s parallel implementation of push relabel[8].
+Much like gluon async [ref1] in distributed Galois [ref3], updates are propagated iteratively until convergence.
+DRONE[ref2] explores a subgraph boundary synchronization scheme over a vertex-cut based partitioning, designed similar to d-Galois but with the additional clause of only boundary nodes having master and mirror nodes and a master-slave architecture where the master node responsible for synchronization and book-keeping.
+DRONE [ref2] claims better performance over gluon[ref1] and gemini [ref4]
+In all three of the above frameworks [ref1] [ref4] [ref2], the propagation of updates to mirror nodes and their synchronization happens in pulses, as defined in Goldberg and Tarjan’s parallel implementation of push relabel[ref8].
 The convergence paradigm represents an inherent universality as it works across shared memory multi-core, distributed memory, as well as GPUs.
 However, for distributed Graph algorithms, every pulse is expensive and bulkier pulses happening less frequently achieving similar results is a faster directive.
-Poly[10] The exploitation of reduction operations in polyhedral analysis are detailed in [10] where multiple possible reduction operations were evaluated to perform 2.21 times faster than control.
+Poly[ref10] The exploitation of reduction operations in polyhedral analysis are detailed in [ref10] where multiple possible reduction operations were evaluated to perform 2.21 times faster than control.
 Semantic knowledge of delay-able statements present in reduction operators allowed scheduling of multi-level iterations without both memory and control barrier delays.
-Cagra[9] The irregular memory access patterns with graph algorithms and their effects when scaled are studied in Cagra[9], that attempts to bring down the cache miss rate by vertex relabeling.
-Cagra [9] does a best effort service to fit the neighborhood of a vertex into the L1/L2 cache of the CPU for shared memory frameworks.
+Cagra[ref9] The irregular memory access patterns with graph algorithms and their effects when scaled are studied in Cagra[ref9], that attempts to bring down the cache miss rate by vertex relabeling.
+Cagra [ref9] does a best effort service to fit the neighborhood of a vertex into the L1/L2 cache of the CPU for shared memory frameworks.
 Distributed memory, especially MPI, would have one process affined to a core, and the vertices managed by a core would be a dividend of the total vertices present in the graph, thus achieving a similar effect with cache misses, in a natural fashion.
-StarPlat[5] a versatile graph DSL provides a distributed backend for generating graph algorithms in openMPI, a more detailed explanation of it is within section III.
-Whereas inclusion of RMA to StarPlat in [6] gave flexibility to code generation, the performance regressed multiple times, small and frequent RMA updates are expensive due to RMA read/update cycle initiation and termination overheads.
-The StarPlat compiler’s[6] distributed backend utilizes a peer-to-peer graph data structure in its header library and generates the openMPI code from the DSL code using StarPlat’s graph AST.
+StarPlat[ref5] a versatile graph DSL provides a distributed backend for generating graph algorithms in openMPI, a more detailed explanation of it is within section III.
+Whereas inclusion of RMA to StarPlat in [ref6] gave flexibility to code generation, the performance regressed multiple times, small and frequent RMA updates are expensive due to RMA read/update cycle initiation and termination overheads.
+The StarPlat compiler’s[ref6] distributed backend utilizes a peer-to-peer graph data structure in its header library and generates the openMPI code from the DSL code using StarPlat’s graph AST.
 In this section we describe essential structures in the StarPlat’s distributed backend.
@@
 Figure 2: Background
-  • Graph Data Structure: The graph partitions, as shown in Figure 2a, in Starplat[5] do not contain ghost or mirror nodes used in gluon-async [1] or DRONE [2].
-Updates are directly and instantaneously propagated, and the pulses used to propagate updates in [1], [2] and [4] are not implemented so that the programmer using StarPlat does writes their code without having any awareness of parallel programming.
+  • Graph Data Structure: The graph partitions, as shown in Figure 2a, in Starplat[ref5] do not contain ghost or mirror nodes used in gluon-async [ref1] or DRONE [ref2].
+Updates are directly and instantaneously propagated, and the pulses used to propagate updates in [ref1], [ref2] and [ref4] are not implemented so that the programmer using StarPlat does writes their code without having any awareness of parallel programming.
 This results in frequent updates that are smaller in size (usually 4 Byte integers).
-    The overhead of initiating and terminating RMA cycles for frequent, small transmissions, is not handled by openMPI[7].
-Therefore, the performance dipped for static Single Source Shortest Paths from [5] to [6] from 3411s to 6331.941s, an increase of 1.856 times.
+    The overhead of initiating and terminating RMA cycles for frequent, small transmissions, is not handled by openMPI[ref7].
+Therefore, the performance dipped for static Single Source Shortest Paths from [ref5] to [ref6] from 3411s to 6331.941s, an increase of 1.856 times.
   • StarPlat’s AST: StarPlat uses its own abstract syntax tree as a form of MLIR.
@@
 The StarPlat compiler’s distributed backend prefers generality in code generation over performance.
-By profiling generated code for Single Source Shortest Paths using gprof [11], perf and hpctoolkit[12], we deduced function calls that were utilizing high number of CPU cycles.
+By profiling generated code for Single Source Shortest Paths using gprof [ref11], perf and hpctoolkit[ref12], we deduced function calls that were utilizing high number of CPU cycles.
 We logged times of these functions and concluded the following phases to be expensive:
@@
     A statement S within a reduction-exclusive statement is also a reduction-exclusive statement.
-The definition of a pulse 3 is an extended, more abstract form of what appears in [8], where delayed and aggregated communication was found beneficial for distributed algorithms.
+The definition of a pulse 3 is an extended, more abstract form of what appears in [ref8], where delayed and aggregated communication was found beneficial for distributed algorithms.
 The lemma 1, allows us to aggregate several pulses 3 and achieve bulkier and less frequent communication, which is ideal.
@@
 To adhere to the peer-to-peer and no mirror node ideology, StarPlat initiates communication with owner nodes and transfers the neighborhood information on demand.
-We noticed through gprof[11], that get_edge calls within neighborhood traversals were taking up a high number of CPU cycles.
+We noticed through gprof[ref11], that get_edge calls within neighborhood traversals were taking up a high number of CPU cycles.
 (a) DSL input.
@@
   • The stack depth for an MPI_Get call is huge.
-OpenMPI[7] does not internally optimize subsequent RMA calls by keeping RMA phases active.
+OpenMPI[ref7] does not internally optimize subsequent RMA calls by keeping RMA phases active.
 Keeping the RMA Window under a shared lock also does not optimize the get calls.